Two planes leave Pittsburgh and Philadelphia at the same time, each going to the other city. One plane flies 25 miles per hour faster than the other. find the air speed of each plane if the cities are 275 miles apart and the planes pass one another after 40 minutes of flying time.
Answer by jorel1380(3719)

2/3x+2/3(x+25)=275
2x+2(x+25)=825
4x+50=825
4x=775
x=193 3/4
One plane flies 193 3/4mph, the other flies at 218 3/4mph.